# A proof of the reverse isoperimetric inequality using a geometric–analytic approach

### v2

@@ I Introduction
This is possible for the AdS case for which we have $\Lambda<0$, which implies that $P>0$ as required on physical grounds. The volume is given as
$\Theta=-\frac{V_{D-2}r_{h}^{D-1}}{(D-1)},$ (2)
where $V_{D-2}$ is the volume of unit sphere and $r_{h}$ is the horizon radius. This leads to the modified first law with a varying cosmological constant as
$\delta M=\frac{\kappa}{8\pi G}\delta A+\frac{\Theta}{8\pi G}\delta\Lambda,$ (3)
where $\kappa$ is the surface gravity.
@@
which is over the interior of the black hole, where we vary the radial coordinate from the singularity at $r=r_{0}$ to the outer horizon at $r=r_{h}$. In fact, the thermodynamic and geometric volumes are the same for all static, spherically symmetric black holes. Therefore, Cvetic et al. restated the conjecture as—“At a fixed geometric volume $V^{\prime}$ , the black hole with the largest entropy is Schwarzschild-AdS”. Furthermore, the geometric volume for the AdS–Schwarzschild black hole evaluates to
@@ II.1 Sherif–Dunsby rigidity and maximal entropy
In the 1+1+21{+}1{+}2 formalism, let $e^{a}$ be the unit vector tangent to the spacelike 22–sheets on $\mathcal{T}$ and $\hat{\theta}\equiv\delta_{\mu}e^{\mu}$ the associated sheet–expansion scalar. In AdS, $R_{ab}e^{a}e^{b}=\Lambda<0$, so the Raychaudhuri equation for $\hat{\theta}$ implies
$\hat{\theta}<0\qquad\text{on }\mathcal{T}.$ (13)
This expresses the gravitational focusing of the spacelike sheet congruence. Since $\hat{\theta}$ equals the logarithmic derivative of the 22–area element along $e^{a}$, the negative sign of (13) corresponds to a contracting conformal rescaling of the intrinsic metric on $\mathcal{T}$. Hence there exists a proper conformal representative
@@
which preserves the scalar curvature, $R(\widetilde{h})=R(h)$. In the 1+1+21{+}1{+}2 decomposition this is precisely the identification made by Sherif and Dunsby [19]: the condition $\varphi<0$ in the conformal transformation is equivalent to a negative sheet expansion on $\mathcal{T}$.
@@
Geometrically, the Sherif–Dunsby rigidity theorem implies that any non–homothetic, scalar–curvature–preserving and volume–preserving deformation of a compact 33–sphere in an Einstein background with negative cosmological constant is dynamically driven back to the round configuration by gravitational focusing. The round $S^{3}$ is therefore the unique stable extremal geometry under these constrained deformations. Thermodynamically, these constraints correspond precisely to fixing the thermodynamic volume $V$ (via volume–preserving deformations) and fixing the pressure $P$ (by holding the background cosmological constant $\Lambda$ fixed). Stability of the round $S^{3}$ thus translates into the statement that it maximizes the entropy at fixed $(P,V)$, with all conserved charges held fixed. In this sense, the round $S^{3}$ represents the maximally entropic spatial slice in the appropriate microcanonical ensemble.
@@ II.2 Effective Functional and Its Variation
where $\gamma_{ij}$ is the induced metric on each constant–$r$ surface, $S_{B}$ denotes the outer York boundary, and $S_{H}$ a small surface of constant proper distance from the bolt. The contribution from $S_{B}$ combines with the explicit Gibbons–Hawking–York term to form the slice functional (22), encapsulating the dependence of the action on the boundary geometry. By contrast, the inner endpoint at the smooth bolt gives a finite, universal residue obtained by evaluating the above expression in the near–horizon geometry $ds^{2}\simeq\rho^{2}\kappa^{2}d\tau^{2}+d\rho^{2}+r_{h}^{2}d\Omega_{D-2}^{2}$, which yields
@@
up to constants. Here, $S$ is the York boundary (D2)(D-2)-sphere. The cosmological constant $\Lambda$ acts as and effective Lagrange multiplier implementing the volume constraint.
@@
where $n^{a}$ is the unit normal and $\phi$ a smooth deformation function on $S$. The first variations of area and volume are given by the Hadamard formulas [^3] The volume functional is defined geometrically as $V[S]=\int_{\text{inside }S}\!\!\sqrt{h}\,d^{3}x$. Since this term arises in the Euclidean action as the quantity conjugate to the cosmological constant $\Lambda$, it corresponds to the thermodynamic volume in extended black–hole thermodynamics. For a static AdS background these two notions coincide, so it is convenient to represent and vary $V$ using the geometric expression $\delta V=\int_{S}\phi\,dA$.
@@
The second variations of area and volume are
@@
Off–shell deformations and Kerr–AdS.
@@
Our second variation calculation establishes that the round sphere is a strict local entropy maximum at fixed thermodynamic volume. The key result is that off–shell deformations of the York boundary functional [^5] The quantity $I_{\rm slice}$ represents the contribution of the outer York boundary to the Euclidean action when evaluated on the regular saddle. It may equivalently be viewed as a geometric functional that captures how the action would change under infinitesimal deformations of the boundary surface at fixed bulk geometry. In this sense, $I_{\rm slice}$ can also be regarded as an “off–shell” functional defined by varying the York surface while keeping the bulk metric fixed, even though in the on–shell decomposition $I_{E}=I_{\rm slice}-\mathcal{A}/4G\,$ it simply represents the boundary part of the total on–shell action. correspond directly to geometric deformations of the horizon area. Specifically, for all spherical harmonic modes $\ell\geq 2$, we find
@@
which shows that the horizon area $\mathcal{A}$ is locally maximized at the round sphere. This result has direct implications for rotating black holes. Since the space of on–shell solutions forms a subset of the off–shell deformation space,
@@
any entropy reduction established for off–shell deformations necessarily holds for the corresponding on–shell solutions. Turning on rotation corresponds precisely to an $\ell=2$ axisymmetric deformation of the round sphere, whose on–shell realization is the Kerr–AdS spacetime. Therefore, Kerr–AdS horizon inherits the entropy reduction property from our general variational analysis. Consequently, along the Kerr–AdS branch the entropy admits the expansion
$S(V,J)\;=\;S(V,0)\;-\;\kappa\,J^{2}\;+\;\mathcal{O}(J^{4})\,,\qquad\kappa>0\,$, (43)
demonstrating that entropy decreases quadratically with angular momentum away from the non–rotating Schwarzschild–AdS solution at fixed thermodynamic volume. Therefore, Kerr–AdS black holes do not maximize entropy at fixed thermodynamic volume—only the spherically symmetric Schwarzschild–AdS solution achieves this maximum. This variational conclusion can also be recovered from a purely thermodynamic argument, which we demonstrate next.
Concavity of $S$ in $J$ at fixed $V$.
We work on the submanifold of thermodynamic state space with fixed thermodynamic volume $V$, and introduce the generalized Helmholtz potential
$\begin{split}&\Phi(T,\Omega;V)\;\equiv\;U-TS-\Omega J\;=\;M-PV-TS-\Omega J,\\ &d\Phi\;=\;-S\,dT\;-\;J\,d\Omega\;-\;P\,dV.\end{split}$ (44)
Here, $T$ is the temperature, $\Omega$ is the angular velocity, $J$ is the angular momentum, $S$ is the entropy, $P$ is the pressure, and U=MPVU=M-PV is the internal energy. The potential $\Phi$ is regarded as a function of the independent variables $(T,\Omega;V)$. A subscript on $\Phi$ indicates a partial derivative with respect to the corresponding variable while holding the others fixed. For example,
$\Phi_{T}\equiv\left(\frac{\partial\Phi}{\partial T}\right)_{\Omega,V},\quad\Phi_{TT}\equiv\frac{\partial^{2}\Phi}{\partial T^{2}}\Big|_{\Omega,V},\quad\Phi_{T\Omega}\equiv\frac{\partial^{2}\Phi}{\partial T\,\partial\Omega}\Big|_{V}.$

### v3

@@ I Introduction
This is possible for the AdS case for which we have $\Lambda<0$, which implies that $P>0$ as required on physical grounds. The conjugate thermodynamic potential is given as
$\Theta=-\frac{(D-2)}{16\pi(D-1)}r_{h}A-\frac{1}{2(D-1)}\sum_{i}a_{i}J_{i},$ (2)
where $a_{i}$ are the rotation parameters, $J_{i}$ are the angular momenta of the black hole, and $A$ is the area of outer horizon. This leads to the modified first law with a varying cosmological constant as
$dM=TdS+\Omega_{i}dJ_{i}+\Theta d\Lambda.$ (3)
@@
which is over the interior of the black hole, where we vary the radial coordinate from the singularity at $r=r_{0}$ to the outer horizon at $r=r_{h}$. In fact, the thermodynamic and geometric volumes are the same for all static, spherically symmetric black holes. For Kerr–AdS, the geometric volume obeys the standard isoperimetric inequality while the thermodynamic volume obeys the reverse isoperimetric inequality. Therefore, Cvetic et al. stated the conjecture as—At a fixed thermodynamic volume $V$ , the black hole with the largest entropy is Schwarzschild-AdS. Furthermore, the geometric volume for the AdS–Schwarzschild black hole evaluates to
@@
Organization of the paper.
The remainder of this article is organized as follows. In Sec. II, we analyze the entropy of AdS–Schwarzschild black holes using the geometric volume, which coincides with the thermodynamic volume in the static case. Using gravitational focusing together with the Sherif–Dunsby conformal rigidity theorem, and independently through a second–variation analysis of the slice functional, we show that the round horizon is the unique entropy maximizer among all hypersurfaces enclosing a fixed thermodynamic volume.
In Sec. III, we extend the discussion to rotating geometries. We first treat Kerr–AdS as an off–shell deformation of the round sphere at fixed thermodynamic volume and show that such a deformation necessarily lowers the entropy relative to the Schwarzschild–AdS case. We then provide an independent thermodynamic argument: since the entropy $S(\mathcal{V},J)$ is strictly concave in $J$ along the thermodynamically stable Kerr–AdS branch, the state with $J=0$ uniquely maximizes entropy at fixed thermodynamic volume $\mathcal{V}$. Together, these two approaches establish that Schwarzschild–AdS is the unique entropy maximizer at fixed thermodynamic volume.
Sec. IV summarizes the main results, delineates the domain of validity of the proof, and outlines several extensions, including modified gravity, non–AdS asymptotics, and quantum or holographic corrections. Appendix A reviews the Sherif–Dunsby rigidity theorem and its relevance to our construction, Appendix B derives the variation of the York boundary term, and Appendix C summarizes the extrinsic geometry of round spheres in Euclidean AdS.
@@ II.1 Sherif–Dunsby rigidity and maximal entropy
In the 1+1+21{+}1{+}2 formalism, let $e^{a}$ denote the unit vector orthogonal to the spacelike 22–sheets on $\mathcal{T}$, and define the associated sheet–expansion scalar by
$\hat{\theta}\equiv\delta_{\mu}e^{\mu}.$ (13)
The quantity $\hat{\theta}$ measures the logarithmic rate of change of the area element of the transverse 22–sheets along the spacelike congruence generated by $e^{a}$.
The evolution of $\hat{\theta}$ is governed by the Raychaudhuri equation for this spacelike congruence. In an Einstein background with negative cosmological constant,
$R_{ab}=\Lambda g_{ab},\qquad\Lambda<0,$ (14)
the Ricci contribution enters with the contracting sign, so that the curvature term drives the expansion toward smaller values along the congruence (gravitational focusing). Once the sheet expansion becomes non–positive at some stage along the flow, the Raychaudhuri equation ensures that the subsequent evolution favors further contraction. We therefore evaluate the rigidity theorem in the regime where
$\hat{\theta}<0,$ (15)
corresponding to a contracting sheet geometry.
Since $\hat{\theta}$ equals the logarithmic derivative of the 22–area element along $e^{a}$, negative sheet expansion corresponds to a contracting conformal rescaling of the intrinsic metric on $\mathcal{T}$. Consequently there exists a conformally related representative
@@
which preserves the scalar curvature, $R(\widetilde{h})=R(h)$. Within the 1+1+21{+}1{+}2 decomposition, this identification between negative sheet expansion and a contracting conformal factor is precisely the relation used by Sherif and Dunsby [23].
@@
Geometrically, the Sherif–Dunsby rigidity theorem, combined with gravitational focusing, implies that any non-homothetic, scalar-curvature-preserving (imposed to isolate pure/genuine shape deformations from changes in the intrinsic curvature scale allowing the geometric argument to probe whether a non-round shape can maximize entropy at fixed thermodynamic volume without introducing additional curvature effects), and volume-preserving deformation of the round compact $S^{3}$ in an Einstein background with negative cosmological constant is not admissible. The round $S^{3}$ is therefore the unique stable configuration under these constrained deformations.
Thermodynamically, these constraints correspond precisely to fixing the thermodynamic volume $V$ (via volume-preserving deformations) and fixing the pressure $P$ (by holding the background cosmological constant $\Lambda$ fixed). Stability of the round $S^{3}$ thus translates into the statement that it maximizes the entropy at fixed $(P,V)$, with all conserved charges held fixed. In this sense, the round $S^{3}$ represents the maximally entropic configuration in the appropriate microcanonical ensemble.
@@ II.2 Effective Functional and Its Variation
where $\gamma_{ij}$ is the induced metric on each constant–$r$ surface, $S_{B}$ denotes the outer York boundary, and $S_{H}$ a small surface of constant proper distance from the bolt. The contribution from $S_{B}$ is regrouped with the non-divergence terms in the Gauss–Codazzi decomposition to yield a volume term, which, when combined with the cosmological constant contribution, forms the bulk volume part of the action. Together with the Gibbons–Hawking–York (GHY) term, this reconstructs the slice functional (24), thereby encapsulating the dependence of the action on the boundary geometry.
By contrast, the inner endpoint at the smooth bolt gives a finite, universal residue obtained by evaluating the above expression in the near–horizon geometry $ds^{2}\simeq\rho^{2}\kappa^{2}d\tau^{2}+d\rho^{2}+r_{h}^{2}d\Omega_{D-2}^{2}$, which yields
@@
up to constants. Here, $S$ is the York boundary (D2)(D-2)-sphere. The cosmological constant $\Lambda$ can be understood as an effective Lagrange multiplier implementing the volume constraint.
@@
where $n^{a}$ is the unit normal and $\phi$ a smooth deformation function on $S$. The first variations of area and volume are given by the Hadamard formulas [^3] The volume functional is defined geometrically as $V[S]=\!\int_{\text{inside }S}\!\!\sqrt{h}\,d^{3}x$. In the static case, this term arises in the Euclidean action as the quantity conjugate to the cosmological constant $\Lambda$, and therefore corresponds to the thermodynamic volume in extended black–hole thermodynamics. Its variation is expressed geometrically by the Hadamard formula, $\delta V=\!\int_{S}\!\phi\,dA$, where $\phi$ denotes the normal deformation of the surface $S$.
@@
The second variation of area and volume are
@@
III Kerr–AdS and the reverse isoperimetric inequality
In this section, we extend our analysis to include rotation and show that it lowers entropy at fixed thermodynamic volume, establishing AdS–Schwarzschild as the unique entropy maximizer.
III.1 Off–shell deformations and Kerr–AdS
Before we begin the analysis for Kerr–AdS case, it is important to note that we compare entropy $S$ of static and rotating case at a fixed thermodynamic volume $\mathcal{V}$ rather than the geometric volume. For the static case, recall that the volume $V[S]$ appears conjugate to the cosmological constant in the Euclidean action and therefore represents the thermodynamic volume, while for the rotating case, they are not the same. More precisely, the slice functional
$I_{\rm slice}[S]=-A[S]-\lambda\,V[S],$ (43)
arises from the Euclidean action of a static black hole. In the static ensemble the Killing vector is
$\xi_{\rm static}=\partial_{\tau},$
which is hypersurface–orthogonal and generates only the Euclidean time circle of period $\beta=1/T$. The boundary data fixed at infinity are therefore exclusively the temperature $T$ (and pressure through $\Lambda$). In this setting the volume conjugate to the pressure is
$\mathcal{V}=\left(\frac{\partial I_{E}}{\partial P}\right)_{T}=\int_{\text{inside}\,S}\sqrt{h}\,d^{3}x=V[S],$ (44)
so the geometric and thermodynamic volumes coincide. Hence Eq. (43) is the correct off–shell functional for studying volume–preserving deformations of the static Schwarzschild–AdS horizon.
For a rotating spacetime the ensemble is different. The relevant Euclidean Killing vector is
$\xi_{\rm rot}=\partial_{\tau}+\Omega\,\partial_{\phi},$
which is not hypersurface–orthogonal and introduces an additional potential $\Omega$ fixed at the boundary. The rotating ensemble is defined by fixing the boundary data $(\beta,\Omega,\Lambda)$, so that the Euclidean time circle is generated by the Killing vector $\xi_{\rm rot}=\partial_{\tau}+\Omega\,\partial_{\phi}$. In this grand–canonical ensemble the generator of Euclidean time translations is $H_{\rm rot}=M-\Omega J$, and the on–shell Euclidean action evaluates to the corresponding Gibbs potential,
$I_{E}^{\rm Kerr}=\beta(M-\Omega J)-S,$ (45)
where $\Omega$ is the angular velocity measured with respect to a non–rotating frame at infinity. One can note that in extended black–hole thermodynamics the mass $M$ is interpreted as the enthalpy, and the thermodynamic volume is defined unambiguously as the quantity conjugate to the pressure in the first law and Smarr relation (equivalently, the Komar/Iyer–Wald volume). In the rotating grand–canonical ensemble $(T,\Omega,P)$ the appropriate thermodynamic potential is the Gibbs free energy $G=M-TS-\Omega J$, which is computed by the Euclidean on–shell action, $I_{E}=\beta G$. Consequently,
$\mathcal{V}=\left(\frac{\partial G}{\partial P}\right)_{T,\Omega}=\frac{1}{\beta}\left(\frac{\partial I_{E}}{\partial P}\right)_{T,\Omega},$ (46)
so the volume extracted from the Euclidean action necessarily coincides with the thermodynamic volume appearing in the Lorentzian first law and Smarr relation. This identification is standard in extended black–hole thermodynamics; see, for example, Dolan [7] and Cheng et al. [3] (see Appendix D for an explicit computation of Dolan’s result in the rotating case). In Cvetič et al. [5], the thermodynamic volume is defined as the quantity conjugate to the pressure in the enthalpy $M$,
$V=\left(\frac{\partial M}{\partial P}\right)_{S,J}.$ (47)
In our Euclidean formulation, the relevant thermodynamic potential in the rotating grand–canonical ensemble is the Gibbs free energy $G=M-TS-\Omega J$, and the volume is defined as
$\mathcal{V}=\left(\frac{\partial G}{\partial P}\right)_{T,\Omega}.$ (48)
Since $G$ is obtained from $M$ by a Legendre transform with respect to $(S,J)$, both definitions yield the same thermodynamic volume.
Taking an off–shell variation at fixed $(\beta,\Omega)$ gives $\delta I_{E}=\beta(\delta M-\Omega\,\delta J)-\delta S$, so consistency of the variational principle requires the boundary functional to contain the term $\beta\Omega J$. Accordingly, integrating out the bulk Einstein–Hilbert–GHY action subject to these grand–canonical boundary conditions yields the rotating slice functional
$I_{\rm slice}^{\rm rot}[S]=-A[S]-\lambda\,V[S]-\beta\,\Omega\,J[S],$ (49)
which is the natural rotating generalization of the static slice functional (43). Here “off–shell” refers to variations of the York slice $S$ at fixed $(\beta,\Omega,\Lambda)$, with the bulk equations of motion otherwise imposed.
It is convenient to recast Eq. (49) in a static–looking form
$I_{\rm slice}^{\rm rot}[S]=-A[S]-\lambda\,\mathcal{V}[S],$ (50)
by defining the “off–shell” thermodynamic volume functional [^5] We emphasize that the York-surface slice functional $I_{\rm slice}[S]$ is an auxiliary off–shell object introduced solely to implement the variational argument. Off shell, there is no unique or canonical notion of “thermodynamic volume” associated with a generic surface $S$. The physical thermodynamic volume entering black–hole chemistry and the reverse isoperimetric inequality is defined on shell via the extended first law and Smarr relation (equivalently via Komar/Iyer–Wald constructions). In the Euclidean formulation, this volume is extracted as $V=(1/\beta)(\partial I_{E}/\partial P)_{T,\Omega}$ in the appropriate ensemble. Our use of $I_{\rm slice}[S]$ therefore does not introduce a new definition of volume, but provides an off–shell variational framework whose on–shell thermodynamics reproduces the standard result.
$\mathcal{V}[S]=V[S]+\frac{\beta\,\Omega\,J[S]}{\lambda}.$ (51)
This shows that, in the rotating case, the geometric volume appearing in the slice functional must be supplemented by rotation–dependent data; the physical thermodynamic volume is defined only on–shell and, for Kerr–AdS, differs from the naive geometric volume by well–known rotation–dependent contributions. As such, if we use the thermodynamic volume $\mathcal{V}$ for comparison, we can represent the rotating slice functional in the form of static slice functional so as to be able to compare the entropy $S$ at the same thermodynamic volume $\mathcal{V}$ exactly as the conjecture demands. Therefore, the comparison of entropy between static and rotating ensembles must be made using the thermodynamic volume and not the geometric volume.
@@ III.1 Off–shell deformations and Kerr–AdS
Our second variation calculation establishes that the round sphere is a strict local entropy maximum at fixed geometric volume (which equals the thermodynamic volume in the static case). The key result is that off–shell deformations of the York boundary functional [^6] The quantity $I_{\rm slice}$ represents the contribution of the outer York boundary to the Euclidean action when evaluated on the regular saddle. It may equivalently be viewed as a geometric functional that captures how the action would change under infinitesimal deformations of the boundary surface at fixed bulk geometry. In this sense, $I_{\rm slice}$ can also be regarded as an “off–shell” functional defined by varying the York surface while keeping the bulk metric fixed, even though in the on–shell decomposition $I_{E}=I_{\rm slice}-\mathcal{A}/4G\,$ it simply represents the boundary part of the total on–shell action. correspond directly to geometric deformations of the horizon area. Specifically, for all spherical harmonic modes $\ell\geq 2$, we find
@@
which shows that the horizon area $\mathcal{A}$ is locally maximized at the round sphere. Moreover, the geometric part of the proof establishes that any finite off–shell deformation away from roundness is not allowed under rigidity and gravitational focusing in AdS space.
These results have direct implications for rotating black holes. Since the space of on–shell solutions forms a subset of the off–shell deformation space,
@@
any entropy reduction established for off–shell deformations necessarily holds for the corresponding on–shell solutions. Turning on rotation corresponds precisely to an $\ell=2$ axisymmetric deformation of the round sphere, whose on–shell realization is the Kerr–AdS spacetime. Therefore, Kerr–AdS horizon inherits the entropy reduction property for all angular momentum $J$ from our general geometric and variational analysis
$S_{\rm Kerr}(\mathcal{V})<S_{\rm Sch}(\mathcal{V}),$ (54)
and do not maximize entropy at fixed thermodynamic volume $\mathcal{V}$—only the spherically symmetric Schwarzschild–AdS solution achieves this maximum.
The fact that Kerr–AdS indeed has strictly lower entropy than Schwarzschild–AdS at fixed thermodynamic volume is confirmed independently in the thermodynamic analysis of $S(\mathcal{V},J)$ in the following subsection, where we work directly in the rotating ensemble.
III.2 Concavity of $S$ in $J$ at fixed thermodynamic volume $\mathcal{V}$
We work on the submanifold of thermodynamic state space with fixed thermodynamic volume $\mathcal{V}$, and introduce the generalized Helmholtz potential
$\begin{split}&\Phi(T,\Omega;V)\;\equiv\;U-TS-\Omega J\;=\;M-P\mathcal{V}-TS-\Omega J,\\ &d\Phi\;=\;-S\,dT\;-\;J\,d\Omega\;-\;P\,d\mathcal{V}.\end{split}$ (55)
Here, $T$ is the temperature, $\Omega$ is the angular velocity, $J$ is the angular momentum, $S$ is the entropy, $P$ is the pressure, and $U=M-P\mathcal{V}$ is the internal energy. The potential $\Phi$ is regarded as a function of the independent variables $(T,\Omega;\mathcal{V})$. A subscript on $\Phi$ indicates a partial derivative with respect to the corresponding variable while holding the others fixed. For example,
$\Phi_{T}\equiv\left(\frac{\partial\Phi}{\partial T}\right)_{\Omega,\mathcal{V}},\quad\Phi_{TT}\equiv\frac{\partial^{2}\Phi}{\partial T^{2}}\Big|_{\Omega,\mathcal{V}},\quad\Phi_{T\Omega}\equiv\frac{\partial^{2}\Phi}{\partial T\,\partial\Omega}\Big|_{\mathcal{V}}.$